1. Find these terms of the sequence {an}, where an =


2 · (−3)n + 5n.


a) a0


b) a1


c) a4


d) a5



1
Expert's answer
2021-12-13T15:28:05-0500

Let us find these terms of the sequence "\{a_n\}," where "a_n = 2 \u00b7 (\u22123)^n + 5n."


a) We get that "a_0 = 2 · (−3)^0 + 5\cdot 0=2."


b) We get that "a_1 = 2 · (−3)^1 + 5\cdot 1=-6+5=-1."


c)We get that "a_4 = 2 · (−3)^4 + 5\cdot 4=162+20=182."


d) We get that "a_5 = 2 · (−3)^5 + 5\cdot 5=-486+25=-461."